OpenStudy (anonymous):

An application of a rational function is T = (AB)/(A+B), which gives the time, T, it takes for two workers to complete a particular task where A & B represent the time it would take for each individual worker to complete the identical task working alone. It takes William 3 hours longer than Timothy to paint a 4x 5 feet bedroom. Respond to the following: Working together, both can complete the job in 2 hours. How long does it take each one to complete the painting job working alone? Provide a similar example, set up the equation and solve the problem. Show you work step by step

hartnn (hartnn):

let it take William A hours to complete the work, and Timothy B hours, It takes William 3 hours longer than Timothy to paint a 4x 5 feet bedroom. ---> A=B+3 ---->(1) Working together, both can complete the job in 2 hours. -----> (AB)/(A+B) = 2 ---->>(2) can you solve those 2 equations simultaneously to get values of A and B ??

OpenStudy (anonymous):

i have no idea where to even start. I have been trying to figure this problem out for at least a hour

hartnn (hartnn):

thats what exactly i gave you, a start...didn't you get how those 2 equation i formed ?

OpenStudy (anonymous):

yes so when they work together there hours are 1hr each

OpenStudy (anonymous):

so A=1 and B=1

hartnn (hartnn):

how you got that ?? no, its incorrect...

OpenStudy (anonymous):

thats my problem I do not know where to get their hours from

hartnn (hartnn):

from (1) you have A =B+3 put this A in equation (2), then you'll have a equation only in terms of B, so that you can find it.

OpenStudy (anonymous):

how do I find the value of A

hartnn (hartnn):

first could you find B ? then A is just B+3

hartnn (hartnn):

(AB)/(A+B) = 2 ((B+3)B) = 2 (B+3+B)

OpenStudy (anonymous):

\[\frac{ (3+b)b }{3+b+b }= \frac{ 3b+b^{2} }{3+2b }\]

hartnn (hartnn):

you'll get a quadratic in B

hartnn (hartnn):

that looks complicated.. try this, ((B+3)B) = 2 (B+3+B)

OpenStudy (anonymous):

\[-b^{2}+b+6\]

OpenStudy (anonymous):

now i need to factor it

hartnn (hartnn):

b^2+3b=4b+6 b^2-b-6=0 this form will be easier, and yes, factor it.

OpenStudy (anonymous):

(b-2)(b-3)

OpenStudy (anonymous):

B=2 or B=3

hartnn (hartnn):

b^2-b-6=0 (b+2)(b-3) = 0 b=-2 , b= 3 but b is time and it cannot be negative, so b=3 got this ?

OpenStudy (anonymous):

yes

hartnn (hartnn):

so, timothy takes 3 hours, and william will take 3+3 =6 hours, righ ?

OpenStudy (anonymous):

yes

hartnn (hartnn):

can you provide similar example on your own ? just change numbers....

hartnn (hartnn):

ask if any doubts

OpenStudy (anonymous):

i think I can now. i think what confused me was the example my teacher put up of this

hartnn (hartnn):

good...